[figure: 9.7 cm]

MAGNETORESITANCE FOR THIN MULTILAYERS

K. Warda^{a}, L. Wojtczak^{a}, D.Baldomir^{b}, M. Pereiro^{b} and J.E. Arias^{b} aSolid State Physics Department, University of Lodz, Poland bDepartamento de Fisica Aplicada Universidad Santiago de Compostela, Spain.

We consider in plane transport properties for trilayer Fe–Cr–Fe within the quasi-classical approach. An influence of geometrical parameters on the values of potential for spin up and down for parallel and antiparallel configuration of trilayer as well as on magnetoresistance are investigated. The different values of potentials, calculated from the density functional theory, are used for calculation of the conductivity in multilayers. We compare the obtained values of the magnetoresitance within the presented model with the results given in the literature.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Krzysztof Warda

Address for correspondence : Department of Solid State Physics University of Łódź Pomorska 149/153 90–236 Łódź

Email address : kwarda@mvii.uni.lodz.pl